OpenStudy (anonymous):

Derivatives... Just want to check my answers... 1. Find y'(x) if xy=x^2+y^2+1 I got y'(x)=y-2x/(2y-x)

OpenStudy (lgbasallote):

i think it's \(\frac{2x-y}{x-2y}\) hmm... SOLUTION: xy' + y = 2x + 2y y' xy' - 2y y' = 2x - y y' (x -2y) = 2x - y was that your solution??

OpenStudy (anonymous):

yeah that's what I got, except I moved the (x-2y) to the other side :). thank you. So I guess I did it right then :p